What is the formula please to find the radius of a sphere when you know the volume? Thanks for your help with this.
Found 2 solutions by mananth, Alan3354:
Answer by mananth(16949)

Volume of Sphere V = 4/3 *pi*r^3
V/3*pi= r^3
cuberoot ( V /3*pi) = r

Answer by Alan3354(69443)

V = 4*pi*r^3/3
r^3 = 3V/(4*pi)
r = cube root(3V/(4pi))
--------------------------
The other tutor let the 4 get away.